OpenStudy (anonymous):

Solve: negative 1 over 2 x + 2 = −x + 7

OpenStudy (lgbasallote):

pls use grouping symbols or that equation tab...

OpenStudy (anonymous):

\[-\frac{1}{2}x+2=-x+7\]??

OpenStudy (istim):

She might be afk.

OpenStudy (anonymous):

afk? loved her in terminator though

OpenStudy (anonymous):

what the answer?